# Motion Capture from Inertial and Vision Sensors

### v1

@@ 6 Conclusion
This paper presents a novel paradigm for consumer-affordable multi-modal motion capture using a monocular camera and sparse IMUs, which offers promising opportunities for personal applications. To support research and applications, we have constructed a large-scale motion capture dataset, called MINIONS, using both inertial and vision sensors. The dataset provides a wide range of representations for human motion, including 3D SMPL models, 2D/3D joints, and action labels, among others. With 5 million frames of 146 fine-grained actions, the dataset is highly scalable for various methods and applications. Moreover, we propose a baseline analysis to learn discriminative representations from IMUs and cameras by extracting supplementary features. The results of experiments underscore the distinct benefits offered by the combination of inertial and vision sensors, highlighting the potential of multi-modal motion capture using a monocular camera and sparse IMUs.

### v3

@@ VI Conclusion
This paper presents a novel paradigm for consumer-affordable multi-modal motion capture using a monocular camera and sparse IMUs, which offers promising opportunities for personal applications. To support research and applications, we have constructed a large-scale motion capture dataset, called MINIONS, using both inertial and vision sensors. The dataset provides a wide range of representations for human motion, including 3D SMPL mesh, 2D/3D joints, and action labels, among others. With 5 million frames of 146 fine-grained actions, the dataset is highly scalable for various applications. Moreover, we propose a comprehensive analysis to learn discriminative representations from IMUs and cameras by extracting supplementary features. The results of experiments underscore the distinct benefits offered by the combination of inertial and vision sensors, highlighting the potential of multi-modal motion capture using a monocular camera and sparse IMUs.